Systems Staff Meeting Notes 4/3/00

Systems Staff Meeting Notes
April 3, 2000

Present: Rich Bennett, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. SIRSI meeting Library Management Meeting will be held in the West staff lounge on Thursday, April 6 at 8:30 AM.
Review Minutes of 4/3/00
  1. O2K SR1 updates - there have been a rash of complaints of it destroying computers. We will continue to go very slowly with it.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. All of the parts for the Terrabyte server have been received.
  2. No distribution list for the Dell computers has been issued yet. Should be out sometime this week.
Networking
  1. We are going to kill off the Unsecured folder on NT2. The FBI software scanning department sent out an email announcement about a 911 worm virus. It propagates through open shares, the worm eats your HD and if it can find a modem - it will dial 911. (the latter aspect is what got the FBI involved). We've made the Unsecure file 'read only' and will kill it 4/10/00. Liaisons have been informed.
  2. Squid Proxy Server problems - late Friday afternoon - usage soared - turns out, people all over the world were using this server. They were probably banned from certain sites on their network and found that they can use our server to get out and bypass the site that are restricted to them. Bill applied filters on Friday - use went way up while users re-tried to gain access. This should not affect campus traffic.
  3. Inventory project - all Library CPU's have been counted, showing approximately 750. Many are not in use; some are still in the original box. Martha should be sending out an email regarding this matter.
  4. Judy Holton - talked to Bobbie - the West wiring project will be delayed for two weeks. The company installing the cable trays, USI, had not even applied for their permit until recently. Judy went with them to insure they applied. Everything is on hold until those cable trays are installed.
  5. Bill came in Sunday to reset the phones to Daylight Savings Time, but they had reset themselves. NT did not put out a message regarding the time change, but Windows did.
  6. Wireless demo - will present one to Resource Services some time next week, once Will returns.
Software Projects Status
  1. Ying has been working on the IR database cube, using a tool called the Seagate Analysis. It creates user manipulability to create a cube of data. It was demoed for Everett & Bobbie on Friday, but it wasn't quite right - the features were a bit unclear. Ying will spend some more time investigating the feature before installing it on the IR database.
  2. Hilton has been working on the Preservation Q&A database. It had some problems, so it was unable to be demoed last week. Cathy Mook is out of the office this week, will attempt to demo it next week.
  3. Library Statistics - Jack Water put out an email to the Health Center & Law libraries with the link for the Library Statistics database. The Health Center called with problems: the database is restricted to the use of IE5 - the Health Center is restricted from using IE. The scripting is a security problem. Bill says that we will get a version of IE5 set up for them to use.
  4. Personnel database changes should be wrapped up today.
Calendar
  1. Will out April 3-7.
  2. Rich out April 5 & 6 - FCLA meetings.
  3. Bill out April 7 & April 14.
  4. Debra out April 18-21.


Back to the Systems Homepage

Last updated April 14, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu